A beam of 5 MeV gamma rays is incident on a block of aluminum that is 72.8 cm thick. What fraction of the gamma rays penetrates the block?
Answer #1

from above figure at incident energy of 10 Mev that is 107  absorption coeffecient is 10-1 = 0.1/cm

therefor 0.9 fraction of incident rays will cross 1 cm of aluminium block..

therefor for 72.8 cm ...fraction of incident rays that will cross it = .9^(72.8) = > 0.00046